OpenStudy (anonymous):

9x^5 − 24x^4 + 16x^3 = 0 Find all real solutions.

OpenStudy (anonymous):

x = 4/3 and 0

OpenStudy (anonymous):

x^3 (-4 + 3 x)^2 = 0\[\left\{\{x\to 0\},\{x\to 0\},\{x\to 0\},\left\{x\to \frac{4}{3}\right\},\left\{x\to \frac{4}{3}\right\}\right\} \]

OpenStudy (anonymous):

x^3(3x-4)^2=0